# Automating Governing Knowledge Commons and Contextual Integrity (GKC-CI) Privacy Policy Annotations with Large Language Models

### v2

@@ Appendix D Baseline Non-fine-tuned performance With N-Shot Learning as N varies
Figure 11. Test set accuracy of non-fine-tuned GPT 3.5TPE with simple prompting and 1, 3 and 5 few shot examples. An accurate annotation is one that is an exact string match.
Appendix E Open-Source Model Performance as Epochs Vary
@@
Appendix F Performance Breakdown For The Three Top Performing Models
Figure 16. A comparison of our top-performing models on positive examples by GKC-CI parameter. GPT3,5_TG_10ep refers the generic GPT-3.5 Turbo model, GPT3,5_TPE_10ep refers to the prompt-engineered version of GPT-3.5 Turbo, and GPT3,5_T2S_10ep refers to the two model GPT-3.5 Turbo system. All models were trained for 10 epochs.
Appendix G Codebook
@@
Appendix H Longitudinal Data
Figure 17. Number of annotated GKC-CI parameters in the privacy policies of 8 prominent companies over time. Policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). Bold grid lines indicate which years’ policies were in the corpus and were annotated. The exact parameter counts displayed in this figure appear in Tables 56 in Appendix H.
@@ Appendix H Longitudinal Data
Appendix I Parameter Variance Data
@@
Appendix J Parameter Density Data
@@
Appendix K Visualizer Demonstration
Figure 18. Side-by-side comparison of Facebook’s 2015 and 2019 policies displayed using the visualizer, showing the GKC-CI annotations in the text. The legend for the color meanings is at the bottom of each GUI window. The comparison highlights how the information flow changes between the policies, with parameter annotations—indicated by text color—emphasizing the increase in specificity in information flows in the 2019 policy.

### v3

@@ Appendix D Baseline Non-fine-tuned performance With N-Shot Learning as N varies
Figure 11. Test-set accuracy of non-fine-tuned GPT 3.5TPE with simple prompting and 1, 3 and 5 few shot examples. An accurate annotation is one that is an exact string match. Note the maximum accuracy of $<$20%.
Appendix E Expanded Model Names
Model Name Model Family Model Size BOS, EOS Tokens Added? Training Objective Epochs
RNN_boseos_10ep N/A: Recurrent Neural Network 67M Yes CLM 10
llama2-7b_boseos_10ep llama2 7B No CLM 10
gpt3,5_tpe_10ep GPT 3.5 Turbo Propetary Propetary CLM 10
gpt3,5_tg_10ep GPT 3.5 Turbo Propetary Propetary CLM 10
gpt3,5_t2s_10ep GPT 3.5 Turbo Propetary Propetary CLM 10
gpt2_xl_boseos_10ep GPT 2 XL (1.5B) Yes CLM 10
gpt2_xl_10ep GPT 2 XL (1.5B) No CLM 10
gpt2_boseos_10ep GPT 2 Base (124M) Yes CLM 10
gpt2_10ep GPT 2 Base (124M) No CLM 10
flan-t5_seq2seq_boseos_1ep Flan-T5 Base (248M) Yes MLM 1
flan-t5_seq2seq_1ep Flan-T5 Base (248M) No MLM 1
flan-t5_large_seq2seq_1ep Flan-T5 Large (783M) No MLM 1
flan-t5_large_seq2seq_boseos_1ep Flan-T5 Large (783M) Yes MLM 1
flan-t5_large_seq2seq_1ep Flan-T5 Large (783M) No MLM 1
Table 4. Model names as they correspond to their specific training/fine-tuning interventions.
Appendix F Open-Source Model Performance as Epochs Vary
@@
Appendix G Performance Breakdown For The Three Top Performing Models
Figure 16. A comparison of our top-performing models on positive examples by GKC-CI parameter. GPT3,5_TG_10ep refers the generic GPT-3.5 Turbo model, GPT3,5_TPE_10ep refers to the prompt-engineered version of GPT-3.5 Turbo, and GPT3,5_T2S_10ep refers to the two model GPT-3.5 Turbo system. All models were trained for 10 epochs. Note that there are only 5 ”Consequence” parameters leading to greater variance on the corresponding bars.
Appendix H Codebook
@@
Appendix I Longitudinal Data
Figure 17. Number of annotated GKC-CI parameters in the privacy policies of 8 prominent companies over time. Policies from the Princeton-Leuven Longitudinal Corpus of Privacy Policies (Amos et al., 2021). Bold grid lines indicate which years’ policies were in the corpus and were annotated. The exact parameter counts displayed in this figure appear in Tables 67 in Appendix I.
@@ Appendix I Longitudinal Data
Appendix J Parameter Variance Data
@@
Appendix K Parameter Density Data
@@
Appendix L Visualizer Demonstration
Figure 18. Side-by-side comparison of Facebook’s 2015 and 2019 policies displayed using the visualizer, showing the GKC-CI annotations in the text. The legend for the color meanings is at the bottom of each GUI window. The comparison highlights information flow changes between the policies, with parameter annotations—indicated by text color—emphasizing an increase in the specificity of information flows in the 2019 policy.
Appendix M Creation of Paragraph-Level Annotations
To create the paragraph-level annotations, we utilize a segmentation approach tailored to capture the structural and contextual integrity of each policy text. Instead of relying on sentence-ending punctuation to define annotation boundaries, we segmented each policy into paragraphs by identifying double newline characters. This method was chosen to reflect natural paragraph divisions in the original documents, preserving the thematic and logical flow intended by the policy authors. Following this segmentation, we proceeded with the annotation creation and benchmarking process using the same methodology used for sentence-level annotations.